OpenStudy (anonymous):

Damien works part-time from home as a data entry professional. He enters data in forms and is paid based on the entry accuracy rate. The table below shows his compensation.

OpenStudy (anonymous):



OpenStudy (anonymous):

Based on his regular performance, the equation below can be used to determine C, the total amount Damien earns a month entering data, where x represents the total number of forms he completes that month. C = 14(0.6x) + 12(0.3x) + 0(0.1x) What is the minimum number of forms he must complete in a month to earn a total of at least $1,080? 90 91 1068 1092

OpenStudy (anonymous):

14(0.6x) + 12(0.3x) + 0(0.1x)=1080 8.4x+3.6x+0=1080 12x=1080 x=1080/12=90 :) and don't say thankew ;)

OpenStudy (anonymous):

How can I not! Thank you, Thank you, Thank you ;)

OpenStudy (anonymous):

okie dokie as ur wish ;) thank u